OpenStudy (nbester98):

Lee works in an insurance call centre where she handles claim enquiries from customers whose surnames start with A to E. Studies have revealed that the volume of calls arriving in Lee’s queue follows a Poisson distribution with an average of 18 per hour. If Lee’s queue receives more than 2 calls in 5 minutes, Lee won’t be able to keep up and customers will abandon their call. a) Determine the probability that Lee’s queue will have abandoned calls during a randomly chosen interval of five minutes. P (At least 2 calls in 5 minutes in queue) F(0) = 18^0*e^-18/0! = 1.5229(10^-8) F(1) = 18

OpenStudy (nbester98):

Just want to know if I'm on the right track here... P (At least 2 calls in 5 minutes in queue) F(0) = 18^0*e^-18/0! = 1.5229(10^-8) F(1) = 18^1*e^-18/1! = 2.7414(10^-7) F(2) = 18^2*e^-18/2! = 2.4673(10^-6) P(>2) = 1 – {1.5229(10^-8) + 2.7414(10^-7) + 2.4673(10^-6)} = .9999972 or 99.99972% chance of abandoned calls in random 5 minute period